OpenStudy (anonymous):

what is 15x-10n-2(3n-40)+2n=180

OpenStudy (anonymous):

15x - 10n - 6n+ 80 +2n = 180 15x -14n =100

OpenStudy (anonymous):

If you meant 15n in place of 15x , then replace x with n and the answer would be 15n - 14n = 100 or n = 100

OpenStudy (anonymous):

15x -10n -6n +80 +2n = 180 15x - 14n = 100